14 - h = 8 please add full explanation, Thank you
Answer by algebrahouse.com(1659)

14 - h = 8
-h = -6 {subtracted 14 from each side of the equation}
h = 6 {divided each side of the equation by -1}

For more help from me, visit: www.algebrahouse.com